OpenStudy (anonymous):

Suppose you buy 3 pizzas for a party. Each pizza is cut into 8 slices. Twenty slices are eaten. How many pizzas were eaten? Write your answer as a mixed number.

OpenStudy (anonymous):

(8x3) - 20?

OpenStudy (tester97):

2 1/2

OpenStudy (anonymous):

thanks @KinzaN @tester97

OpenStudy (tester97):

welcome :)

OpenStudy (♪chibiterasu):

8 • 3 = 24 24/20 = 1.2

OpenStudy (♪chibiterasu):

Wait I did that wrong.

OpenStudy (anonymous):

Write 4 and 5 eighths as an improper fraction and as a mixed number.

OpenStudy (anonymous):

If x is number of slices. 3x = number of slices available. Each pizza has 8 slices, so 3*8 = total slices at the party. 20 is total number eaten 20 slices out of (divided by) the total number available = amount of the pizzas eaten. y = 20/3x

OpenStudy (♪chibiterasu):

8 8 8 8 + 8 + 4 2.5 yep.

OpenStudy (tester97):

37/8 4 5/8

OpenStudy (♪chibiterasu):

tester, y u beat me to questions?!?! ;-;